Testing the goodness of fit between the data and the Hardy Weinberg equilibrium model generated expectations.

4.1 In a species of bird, feather color is controlled by genes at a single locus, with the red feather allele dominant to the yellow feather allele. A population has 22 red and 14 yellow birds, with 9 of the red birds having a homozygous dominant genotype. Is this population in equilibrium?

Calculate p and q from the number of individuals of each genotype:

p = ____

q = ____

Calculate the expected frequency of each genotype if the population is in equilibrium:

_____ = Frequency of homozygous dominant individuals

_____ = Frequency of heterozygous individuals

_____ = Frequency of homozygous recessive individuals

Calculate the expected number of individuals of each genotype in a population of 36 birds if the gene is in equilibrium:

_____   = Number of homozygous dominant individuals

_____   = Number of heterozygous individuals

_____ = Number of homozygous recessive individuals

Test how well your data fits the expected values from the equilibrium model:

_____ = Chi-square test statistic

_____ = P value

_____ (y/n) in equilibrium?

4.2 In a species of mouse, tail length is controlled by genes at a single locus, with the long tail allele dominant to the short tail allele. A population has 49 long tail and 25 short tail mice, with 22 of the long tail mice having a homozygous dominant genotype. Is this population in equilibrium?

Calculate p and q from the number of individuals of each genotype:

p = _____

q = _____

Calculate the expected frequency of each genotype if the population is in equilibrium:

_____   = Frequency of homozygous dominant individuals

_____   = Frequency of heterozygous individuals

_____   = Frequency of homozygous recessive individuals

Calculate the expected number of individuals of each genotype in a population of 74 mice if the gene is in equilibrium:

_____   = Number of homozygous dominant individuals

_____   = Number of heterozygous individuals

_____   = Number of homozygous recessive individuals

Test how well your data fits the expected values from the equilibrium model:

_____ Chi-square test statistic

______ P value

_____ (y/n) in equilibrium?


4.3 In humans, the hitchhiker’s thumb trait is controlled by genes at a single locus, with the non-hitchhiker’s thumb allele dominant to the hitchhiker’s thumb allele. A population has 46 people that do not have the hitchhiker’s thumb and 21 that do. Of the humans without a hitchhiker’s thumb, 19 have a homozygous dominant genotype. Is this population in equilibrium?

Calculate p and q from the number of individuals of each genotype:

p = _____

q = _____

Calculate the expected frequency of each genotype if the population is in equilibrium:

_____   = Frequency of homozygous dominant individuals

_____   = Frequency of heterozygous individuals

_____   = Frequency of homozygous recessive individuals

Calculate the expected number of individuals of each genotype in a population of 67 humans if the gene is in equilibrium:

_____   = Number of homozygous dominant individuals

_____   = Number of heterozygous individuals

_____   = Number of homozygous recessive individuals

Test how well your data fits the expected values from the equilibrium model:

_____   Chi-square test statistic

______ P value

_____ (y/n) in equilibrium?


4.4 In a certain species of prickly pear, having straight or curved spines is a trait controlled by genes at a single locus, with the straight spine allele dominant to the curved spine allele. A population of prickly pears has 37 individuals with straight spines and 42 individuals with curved spines. Of the prickly pears with straight spines, 12 have a homozygous dominant genotype. Is this population in equilibrium?

Calculate p and q from the number of individuals of each genotype:

p = _____

q = _____

Calculate the expected frequency of each genotype if the population is in equilibrium:

_____   = Frequency of homozygous dominant individuals

_____   = Frequency of heterozygous individuals

_____   = Frequency of homozygous recessive individuals

Calculate the expected number of individuals of each genotype in a population of 79 prickly pear cacti if the gene is in equilibrium:

_____   = Number of homozygous dominant individuals

_____   = Number of heterozygous individuals

_____   = Number of homozygous recessive individuals

Test how well your data fits the expected values from the equilibrium model:

_____   Chi-square test statistic

______ P value

_____ (y/n) in equilibrium?
Answer #1

4.1) No of homozygous dominant individuals = 9

Total no of individuals = 36

Hence p2 = PP = 9/36 = 0.25

Therefore, p = 0.5

q2 = 14/36 = 0.38

q = 0.62

Hence p + q is not equal to 1 in this case hence the population is not in hardy weinberg equilibrium.

frequency of genotype AA = PP = 9/36 = 0.25

frequency of genotype aa = q2 = 14/36 = 0.38

frequency of genotype Aa= 13/36= 0.36

If the gene is in equilibrium,

No of AA = 9, No of aa = 14, No of Aa = 13

p2 + 2pq + q2 = 1

chi square test static = (O-E)2 / E = 0.00038

This is not in hardy weinberg equilibrium

4.2) No of homozygous dominant = 22

total no = 74

PP = 22/74 = 0.29

p = 0.54

q = 0.57

Not in hardy weinberg equlibrium

Frequency of AA = 22/74 = 0.29 , Aa= 27/74= 0.36 , aa = 25/74 = 0.33

No of AA = 22 , Aa= 27 and aa= 25

Chi square = 0.00129

4.3) PP = 19/ 67 = 0.28 =p = 0.52

pp= 21/67= 0.31 =q= 0.55

p= 0.52, q = 0.55, hence not in hardy weinberg equilibrium since p=+ q is not equal to 1.

Frequency of AA= 0.28, Aa = 27/67= 0.40 , aa= 0.31

No of AA= 19, Aa= 27, aa = 21

chi square= 0.00034

4.4) PP= 12/ 79 = 0.15 = p= 0.38

pp= 42/79= 0.53= q= 0.72

p= 0.38, q= 0.72

hence not in hardy weinberg equilibrium

Frequency of AA= 0.15, Aa= 25/79= 0.31 , aa= 0.53

No of AA= 12, Aa= 25, aa= 42

chi square= 0.0003

This is not in hardy weinberg equilibrium.